# Asymmetry Analysis of Bilateral Shapes

### v1

@@ 3 Hypothesis tests
Consider a dataset of $N$ basis registered configurations observed on two groups of individuals, where the first $N_{1}$ configurations belong to one group and the last $N_{2}$ configurations belong to the other group, $N=N_{1}+N_{2}$, where it is thought that the first group of configurations might be less asymmetric than the second. One way to think about this question is to construct a one-sided hypothesis test of
$\begin{split}&H_{0}:\text{ the two groups have the same distribution of asymmetry vs.}\\ &H_{1}:\text{ Group 1 is less asymmetric than Group 2.}\end{split}$ (3.1)
We will give two general approaches to construct a test statistic: (a) combine-then-compare, and (b) compare-then-combine. These two approaches will be explored in the next subsections.
To specify the null and alternative hypotheses explicitly, some notations are needed. Suppose the data take the form of $N$ registered configurations
@@
$\begin{split}&H_{0}:\text{ $\bm{a}_{n}\sim F$ i.i.d. for all $n$ vs.}\\ &H_{1}:\text{ $\bm{a}_{n}\sim F_{1}$ i.i.d. for $n=1,\dots,N_{1}$, $\bm{a}_{n}\sim F_{2}$ i.i.d. for $n=N_{1}+1,\dots,N$,}\end{split}$ (3.3)
where $F$, $F_{1}$, $F_{2}$ are multivariate cumulative density functions (c.d.f.s) and $F_{1}$ is stochastically smaller than $F_{2}$ (that is, $F_{1}(\bm{a})\geq F_{2}(\bm{a})$ for all $\bm{a}$ with at least one strict inequality), see, for example Stoyan (1983). In other words, the distribution of $F_{2}$ is shifted towards right comparing with $F_{1}$, which indicates more asymmetries.
We concentrate in this section on t-test and Mann-Whitney U test. In Section 4, we further discuss some alternative univariate tests as well as multivariate tests in relation to the smile data.
@@ 3.1 Testing strategy: combine-then-compare
say. Here the function $\phi(\bm{a})$ is assumed to be monotonically increasing in the sense that it is monotonically increasing in each component $a_{j}$ when the other elements are held fixed.
@@
    greater weights near the midplane,
@@
For example, in cases (b) and (c) a weight might depend on $|X[k_{R},1]-X[k_{L},1]|$ for a landmark pair, with solos treated separately. The choices of weights in general depend on prior belief of the extent of asymmetry information carried by each landmark, for example, if it is known a priori that landmarks far away from $\mathcal{P}$ can carry important asymmetry information than landmarks near $\mathcal{P}$, then more weights should be assigned to the farther landmarks. For smile example, this information is available so greater weights will be given near the midplane, see Section 4.
The choice of function $\psi$ governs the relative influence of different magnitudes of an elementary feature. For example, the quadratic function $\psi(a)=a^{2}$ is more sensitive to outlying features than the linear function $\psi(a)=a$.
We write below explicitly for two such scores with equal weights and give them names $L_{1}$ and $L_{2}$ statistics respectively.
[…]
In the above equations, we use the two expressions for elements in $\bm{d}$ and $\bm{a}$ introduced in Section 2.2.
Bock and Bowman (2006) have proposed an asymmetry score with weights $w_{j}=1$ for landmark pairs and weights $w_{j}=2$ for solos and $\psi(a_{j})=a_{j}^{2}$ in equation (3.5), namely,
$\sum_{(k_{L},k_{R})}\sum_{m=1}^{M}d[(k_{L},k_{R}),m]^{2}+2\sum_{k_{S}}d[(k_{S})]^{2}.$ (3.8)
For our work, we have selected an equal weights for $\phi_{L_{1}}$ to have some similarity to the asymmetry score of Patel et al. (2023); their score is given by
$\phi_{L_{1}}^{*}(\bm{a})=\sum_{(k_{L},k_{R})}d^{*}[(k_{L},k_{R})]+\sum_{k_{S}}|d[(k_{S})]|,$ (3.9)
with weights equal to 1 for both pairs and solos. The $d^{*}[(k_{L},k_{R})]$ is defined in equation (2.4). Also, we can define a similar score as follows using $L_{2}$ distances
$\phi_{L_{2}}^{*}(\bm{a})=\sum_{(k_{L},k_{R})}d^{*}[(k_{L},k_{R})]^{2}+\sum_{k_{S}}|d[(k_{S})]|^{2}.$ (3.10)
It can be shown using (2.4) that this equation (3.10) reduces to (3.7). That is, we have
$\phi^{*}_{L_{2}}(\bm{a})=\phi_{L_{2}}(\bm{a}).$
Note that, we can generalize $\phi^{*}(\bm{a})$ in the same way as $\phi(\bm{a})$ with weights.
Example in Section 2.3 continued. Recall the asymmetric square $X_{2}$ defined in equation (2.6). The configuration is shown in Figure 2. The landmark elementary feature, $d^{*}[(1,3)]$, defined in equation (2.4) is computed and reported in Table 2. The composite asymmetry scores $\phi_{L_{1}}(\bm{a})$ (3.6), scaled $\phi^{*}_{L_{1}}(\bm{a})$ (3.9) (divided by the number of landmarks) and $\phi_{L_{2}}(\bm{a})$ (3.7) are computed on $X_{2}$ for illustration. The results are shown in Table 2. Note that the values of
$d^{*}[(1,3)]=\sqrt{(X_{2}[1,1]+X_{2}[3,1])^{2}+(X_{2}[1,2]-X_{2}[3,2])^{2}}=0.18$
@@
Let
@@
denote the expectation and variance for a random variable $U_{g}\sim F_{g}$, where $g=1,2$ labels one of the two groups. A simpler version of hypotheses in (3.1) is
@@
    the standard two-sample t-test, which assumes a common variance for the two groups ($\sigma_{1}^{2}=\sigma_{2}^{2}$);
@@
    the Mann-Whitney U test, also known as the Wilcoxon rank sum test (Mann and Whitney, 1947).
The Mann-Whitney U test is nonparametric so that the test results are invariant under monotonic transformations of the $u_{n}$. In case (a), the significance of the test statistic under $H_{0}$ can be computed either analytically (assuming normality) or using a bootstrap approximation. In case (b) the significance can be derived from combinatorial arguments. Mann-Whitney U test is considered where the normality assumption on composite asymmetry scores $u_{n}=\phi(\bm{a}_{n})$ may fail to hold.
An advantage of the combine-then-compare approach is that it enables us to accumulate evidence from different features. If there is a small amount of asymmetry on many different elementary features, then the composite score will have a large value. Further, this approach enables the two groups to be compared visually with stem and leaf plots and in particular, any overlap between the two groups can be easily assessed. Applications of this approach are in Section 4
@@ 3.2 Testing strategy: compare-then-combine
In this approach, separate test statistics are constructed for each feature in the elementary feature vectors. The most extreme of these separate test statistics is then used as an overall test statistic. Before we give the details, the hypotheses in equation (3.1) need to be formulated in the union intersection test style (as explained in Section 1) as following:
@@
For each choice $j$ of an unsigned elementary feature, construct a statistic to compare the two groups, for example, a two-sample t-statistic or a Mann-Whitney U statistic. Denote the resulting statistics by $v_{j},\ j=1,\ldots,J$. For example, if a t-statistic is used, then
@@
where $\bar{a}_{j}^{(g)},\ g=1,2$ are the sample means of the $j$th unsigned elementary feature in the two groups and $s^{2}_{j}$ is the pooled within-group variance.
@@
An overall UIT statistic to test $H_{0}$ vs $H_{1}$ in equation (3.14) can be defined by taking the maximum of these separate statistics (Boyett and Shuster, 1977),
$V=\max_{j=1,\ldots,J}v_{j}.$ (3.16)
@@
The critical points of $V$ under the null hypothesis can be computed by the bootstrap. Note that the landmarks are correlated with each other, so the $d_{j}$ are not independent for $j=1,\ldots,J$ here, so we cannot derive the distribution of $V$ theoretically. Hence, bootstrap is used for estimating the critical points of $V$.
@@
If there exists $v_{j}$ which exceeds the critical values of $V$, then we not only know that $H_{0}$ is rejected, but also know that the landmark corresponding to such $v_{j}$ is also important.
@@
Our main aim is somewhat different from Boyett and Shuster (1977): we aim at selecting important landmarks while Boyett and Shuster (1977), attempt to estimate a $p$-value. The bootstrap procedure used by us is also different from their who have used:
  1. 1.
    A set $E$ containing all permutations of the dataset is constructed.
  2. 2.
@@
    Sample with replacement from this set $E$, i.e. sample permutations of the data. The permuted samples constitute the bootstrap resampled dataset.
  3. 3.
    Test statistic $V$ is computed on the resampled dataset.
On the other hand, we directly sample from the whole dataset as following:
  1. 1.
@@
    Sample $N$ samples with replacement from $\{\bm{a}_{n}\}_{n=1}^{N}$.
  2. 2.
    Compute $v_{j}$ and $V$ on resampled dataset.
@@
An advantage of this compare-then-combine approach versus combine-and-compare is that if the null hypothesis is rejected, then simultaneous confidence intervals can be constructed to identify those features on which the two groups significantly differ. The selection procedure in spirit is similar to Tukey’s method (Tukey, 1949) in ANOVA though in ANOVA a separate test statistic is used for each pairwise difference between main effects. Here, a separate test statistic for each $j$ is more appropriate since all features in unsigned elementary feature vector $\bm{a}$ have different biology. An illustration is given below for the smile data.
4 Analyses of the Smile Data
We now give smile data taken from Patel et al. (2023) which have been pre-registered. The data is collected using the Di4D system. When a subject’s face is placed in front of the camera, the $x$-axis is positioned horizontally on the face (positive direction is left), the $y$-axis is placed vertically on the face (positive direction is upward), whereas the $z$-axis is positioned in the direction inwards and outwards of the face (positive direction is outward), see Figure 1.
The coordinates given in Section 2.1 can be related to these coordinate axes as follows:
coordinate 1: left-right $x$
coordinate 2: down-up $y$
coordinate 3: back-front $z$.

### v2

@@ 2.3 Example 1
Therefore, there is one landmark pair $(k_{L},k_{R})=(1,3)$ and two solos, $k_{S}=2,4$, so $K_{P}=1$ and $K_{S}=2$. Further, suppose both configurations are treated as basis registered with $\mathcal{P}$ normal to the first coordinate axis, i.e. $\mathcal{P}$ is the $y$-axis.
The configurations $X_{1}$ and $X_{2}$ are shown in the left and right figures in the top row of Figure 2 respectively, whereas the bottom row of Figure 2 show their corresponding reflections, where the left is for $X_{1}$ and $X_{1}^{\text{(refl)}}$ and right is for $X_{2}$ and $X_{2}^{\text{(refl)}}$. Note that $X_{1}$ is bilaterally symmetric, with $\bm{a}=\bm{0}$. Table 1 gives the $J_{\text{basis}}=4$ elements of the vector $\bm{a}$ for $X_{2}$. The two elements for the landmark pair are listed first, and the two elements for the solos listed last, with
$\bm{a}=(0.04,\ 0.18,\ 0.28,\ 0.31)^{T}$
for $X_{2}$ . It can be seen visually in Figure 2 that for $X_{2}$, landmarks 2 and 4 deviate more from the $y$-axis than the landmark pair (1, 3) differs from symmetry. Hence the last two components of $\bm{a}$ are larger than the first two components.
@@ 3 Hypothesis tests
Consider a dataset of $N$ basis registered configurations observed on two groups of individuals, where the first $N_{1}$ configurations belong to one group (Group 1) and the last $N_{2}$ configurations belong to the other group (Group 2), $N=N_{1}+N_{2}$, where it is thought that the first group of configurations might be more symmetric than the second. One way to think about this question is to construct a one-sided hypothesis test of
[…]
For the paired data, we change the Group 1 and Group 2 in the above hypotheses to experimental and control groups respectively, where same subjects are contained in the two paired groups.
We will give two general approaches to construct a test statistic in order to compare the symmetries between the two groups: (a) combine-then-compare, where univariate test methods are used, and (b) compare-then-combine, where we form the hypotheses under union intersection test framework. These two approaches will be explored in the next subsections.
Suppose the data take the form of $N$ registered configurations
@@
[…]
where $F$, $F_{1}$, $F_{2}$ are multivariate cumulative density functions (c.d.f.s) and $F_{1}$ is stochastically smaller than $F_{2}$ (that is, $F_{1}(\bm{a})\geq F_{2}(\bm{a})$ for all $\bm{a}$ with at least one strict inequality, see, for example Stoyan (1983)). In other words, the distribution of $F_{2}$ is shifted towards right comparing with $F_{1}$, which indicates more asymmetries. Under the paired data setting, $\bm{a}_{n}$ for $n=1,\dots,N_{1}$ represent the experimental data and $\bm{a}_{n}$ for $n=N_{1}+1,\dots,N$ represent the control data, where $N=2N_{1}$.
We concentrate in this section on t-test and Mann-Whitney U test. In Section 5, we further discuss some alternative univariate tests as well as multivariate tests in relation to the application of Smile Data 1.
@@ 3.1 Testing strategy: combine-then-compare
say. Here the function $\phi(\bm{a})$ is assumed to be monotonically increasing in each component $a_{j}$ when the other elements are held fixed.
@@
    greater weights near the midplane or away from the midplane, or
@@
For example, in case (b), a weight might depend on $|X[k_{R},1]-X[k_{L},1]|$ for a landmark pair, with solos treated separately. The choices of weights in general depend on prior belief of the extent of asymmetry information carried by each landmark, for example, if it is known a priori that landmarks near $\mathcal{P}$ can carry more important asymmetry information than landmarks far away from $\mathcal{P}$, then greater weights should be assigned to the nearer landmarks. For Smile Data 1, this information is available so greater weights will be given near the midplane, see Section 5.
Another way to give some elementary features greater prominence than others is through the choice of the functions $\psi$. For example, the quadratic function $\psi(a)=a^{2}$ is more sensitive to outlying features than the linear function $\psi(a)=a$.
We write below explicitly two such scores with equal weights and give them names $L_{1}$ and $L_{2}$ statistics respectively.
$\text{$L_{1}$ statistic: }\phi_{L_{1}}(\bm{a})=\sum_{j=1}^{J}a_{j}=\sum_{(k_{L},k_{R})}\sum_{m=1}^{M}|d[(k_{L},k_{R}),m]|+\sum_{k_{S}}|d[(k_{S})]|,$ (3.6)
$\text{$L_{2}$ statistic: }\phi_{L_{2}}(\bm{a})=\sum_{j=1}^{J}a_{j}^{2}=\sum_{(k_{L},k_{R})}\sum_{m=1}^{M}d[(k_{L},k_{R}),m]^{2}+\sum_{k_{S}}d[(k_{S})]^{2},$ (3.7)
where the $d$’s are given in (2.1)-(2.3) and $\bm{a}$, $a_{j}$ are given in (2.5).
We note that Bock and Bowman (2006) have proposed an asymmetry score which is proportional to $\phi_{L_{2}}$, that is
$\frac{2}{K}\phi_{L_{2}}(\bm{a}).$ (3.8)
Further, Patel et al. (2023) uses the score given by
$\phi_{L_{1}}^{*}(\bm{a})=\sum_{(k_{L},k_{R})}d^{*}[(k_{L},k_{R})]+\sum_{k_{S}}d^{*}[(k_{S})],$ (3.9)
The $d^{*}[(k_{L},k_{R})]$ and $d^{*}[(k_{S})]$ are defined in equation (2.4). Also, by squaring each element in the two sums in (3.9) and using (2.4), we recover $\phi_{L_{2}}$:
$\phi_{L_{2}}^{*}(\bm{a})=\sum_{(k_{L},k_{R})}d^{*}[(k_{L},k_{R})]^{2}+\sum_{k_{S}}|d^{*}[(k_{S})]|^{2}=\phi_{L_{2}}(\bm{a}).$ (3.10)
Note that we can generalize $\phi^{*}(\bm{a})$ in the same way as $\phi(\bm{a})$ with weights. Instead of considering asymmetries with respect to each landmark as in (3.9), (3.6) and (3.7) view all features in $\bm{a}$ as a whole and take the $L_{1}$ and $L_{2}$ norms. For the subjects which are more symmetric, their corresponding $\bm{a}_{n}$ should be close to the origin. Hence, the $L_{1}$ and $L_{2}$ distances between $\bm{a}_{n}$ with $\bm{0}\in\mathbb{R}^{J}$ are expected to be smaller for control subjects.
Example 1 continued. Recall the asymmetric square $X_{2}$ defined in equation (2.6). The configuration is shown in Figure 2. The landmark elementary feature, $d^{*}[(1,3)]$, defined in equation (2.4) is computed and reported in Table 2. The composite asymmetry scores $\phi_{L_{1}}(\bm{a})$ (3.6), scaled $\phi^{*}_{L_{1}}(\bm{a})$ (3.9) (divided by the number of landmarks) and $\phi_{L_{2}}(\bm{a})$ (3.7) are computed on $X_{2}$ for illustration. The results are shown in Table 2. Note that the values of
$\begin{split}d^{*}[(1,3)]=\sqrt{(X_{2}[1,1]+X_{2}[3,1])^{2}+(X_{2}[1,2]-X_{2}[3,2])^{2}}=0.18\end{split}$
@@
Suppose $u_{n}$ for $n=1,\ldots,N_{1}$ and $n=N_{1}+1,\ldots,N$ are realizations from random variables $U_{1}$ and $U_{2}$ respectively, and let $g=1,2$ denote the control and cleft groups. Let
@@
denote the expectation and variance for a random variable $U_{g}\sim F_{g}$ (should not be confused with the distribution of the vector $\bm{a}$ in Section 3). A simpler version of hypotheses in (3.1) is
@@
    Parametric test: the standard two-sample t-test, which assumes $\sigma_{1}^{2}=\sigma_{2}^{2}$;
@@
    Non-parametric test: the Mann-Whitney U test, also known as the Wilcoxon rank sum test (Mann and Whitney, 1947).
In case (a), the significance of the test statistic under $H_{0}$ can be computed either analytically (assuming normality) or using a bootstrap approximation. In case (b) the significance can be derived from combinatorial arguments. Mann-Whitney U test is considered where the normality assumption on composite asymmetry scores $u_{n}=\phi(\bm{a}_{n})$ may fail to hold.
Paired data The test strategies described above can be applied to the paired data by making necessary modification, namely we need to use the paired test in place of the two-sample test for both the parametric and non-parametric cases. Specifically, paired t-test and Wilcoxon signed-rank test are used for the parametric and non-parametric cases respectively.
An advantage of the combine-then-compare approach is that it enables us to accumulate evidence from different features, so if there is a small amount of asymmetry on many different elementary features, then the composite score will have a large value. Further, this approach enables the two groups to be compared visually with stem and leaf plots and in particular, any overlap between the two groups can be easily assessed. Applications of this approach are in Section 5.2.1, 6.2.1 and 6.3.1.
@@ 3.2 Testing strategy: compare-then-combine
In this approach, separate test statistics are constructed for each feature in the elementary feature vectors. The most extreme of these separate test statistics is then used as an overall test statistic. Before we give the details, the hypotheses in equation (3.1) need to be formulated in the Union Intersection Test (UIT) style:
@@
For each choice $j$ of an unsigned elementary feature, we construct a statistic to compare the two groups, for example, a two-sample t-statistic or a Mann-Whitney U statistic. Denote the resulting statistics by $v_{j},\ j=1,\ldots,J$. For example, if a t-statistic is used, then
@@
where $\bar{a}_{j}^{(g)},\ g=1,2$ are the sample means of the $j$th unsigned elementary feature in the two groups and $s^{2}_{j}$ is the corresponding pooled within-group variance.
@@
Following Boyett and Shuster (1977), an overall UIT statistic to test $H_{0}$ vs $H_{1}$ in equation (3.14) can be defined by taking the maximum of these separate statistics,
$V=\max_{j=1,\ldots,J}v_{j},$ (3.16)
@@
and Boyett and Shuster (1977) compute the critical points of $V$ under the null hypothesis using bootstrap. The landmarks are correlated with each other, so the $d_{j}$ are not independent for $j=1,\ldots,J$ here, so we cannot derive the distribution of $V$ theoretically. Hence, the bootstrap has been used for estimating the critical points of $V$.
@@
If there exists $v_{j}$ which exceeds the critical value of $V$ at a pre-determined significance level, then we not only know that the corresponding $H_{0}^{j}$ is rejected (hence the overall $H_{0}$ in (3.14) is also rejected), but also know that the landmark corresponding to such $v_{j}$ is important.
@@
Our main aim is somewhat different from Boyett and Shuster (1977): we aim at discovering important landmarks while Boyett and Shuster (1977) focuses on estimating a $p$-value. Also, the two bootstrap procedures are slightly different. Their procedure is as follows:
  1. (a)
    Construct a set $E_{n}$ containing all random samples of size $n$ taken without replacement from origin dataset $\{\bm{a}_{i}\}_{i=1}^{N}$.
  2. (b)
@@
    Sample with replacement $B$ times from this set $E_{n}$ to form the bootstrap resampled dataset.
  3. (c)
    Compute the test statistic $\{V^{(1)},\dots,V^{(B)}\}$ on the resampled dataset.
  4. (d)
    Compare the test statistic $V$ computed on first $n$ observed data $\{\bm{a}_{i}\}_{i=1}^{n}$ to $\{V^{(1)},\dots,V^{(B)}\}$.
The procedure given in Boyett and Shuster (1977) does not perform bootstrap directly on the data, and only part of the data is used in either $E_{n}$ or the observed statistic, which is probably to reduce the computational budget. In contrast, since our data size is not large and the computational budget is far more than enough, we directly sample with replacement from the whole dataset with $n=N$ and use all data to compute the observed statistic. Using all data is very likely to improve the final result. Our procedure is as follows:
  1. (a)
@@
    Sample $N$ samples with replacement from $\{\bm{a}_{i}\}_{i=1}^{N}$ and repeat this step $B$ times.
  2. (b)
    Compute $\{V^{(1)},\dots,V^{(B)}\}$ on resampled dataset.
  3. (c)
    Compare the test statistic $V$ computed on all observed data $\{\bm{a}_{i}\}_{i=1}^{N}$ to $\{V^{(1)},\dots,V^{(B)}\}$.
@@
For compare-then-combine approach, we use a similar selection procedure to Tukey’s method (Tukey, 1949) in ANOVA though in ANOVA a separate test statistic is used for each pairwise difference between main effects. Here, a separate test statistic for each $j$ is more appropriate since the different features in the unsigned elementary feature vector $\bm{a}$ measure different features of the face. An illustration is given below in Section 5.2.2 and 6.3.2.
Paired data The approach given in this section can be extended to the paired data scenario with the following changes applied to the bootstrap procedure.
  1. (a)
    Compute $a^{(2)}_{nj}-a^{(1)}_{nj}$ for each subject $n$ and each feature $j$, where $a^{(1)}_{nj}$ and $a^{(2)}_{nj}$ are the feature $j$ of vectors $\bm{a}_{n}^{(1)}$ and $\bm{a}_{n}^{(2)}$ for subject $n$ from Group 1 and 2 respectively.
  2. (b)
    Assign +1+1 or 1-1 to each $\bm{a}^{(2)}_{n}-\bm{a}^{(1)}_{n}$ randomly with equal probability, which forms the permuted sample. Repeat this step $B$ times.
  3. (c)
    Compute $t$-values on the permuted sample and record the largest $t$-value.
An illustration of this compare-then-combine approach is given below for the Smile Data 2 in Section 6.2.2.
4 Meta-Analysis Methods
We can consider the meta-analysis, which carries out simultaneous inference and also belongs to the compare-then-combine approach category. In particular in meta-analysis, we can obtain an overall $p$-value by combining the $p$-values from separate tests. We now give an overview for meta-analysis.
Suppose we have $N$ basis registered configurations, where the first $N_{1}$ configurations are from Group 1 and the rest $N_{2}$ configurations are from Group 2. Absolute elementary feature vectors $\bm{a}_{1},\dots,\bm{a}_{N}\in\mathbb{R}^{J}$ defined in equation (2.5) can be constructed on each configuration. Let $\bm{a}^{(g)}$ be the random vector associated to the population of group $g$ and let $a_{j}^{(g)}$ denote feature $j$ of vector $\bm{a}^{(g)}$, where $g=1,2$ for Group 1 and Group 2 respectively and $j=1,\dots,J$.
Consider the $J$ individual tests with corresponding hypotheses
$\begin{split}H_{0}^{j}:a_{j}^{(1)}=a_{j}^{(2)},\ \ H_{1,L}^{j}:a_{j}^{(1)}<a_{j}^{(2)},\ \ H_{1,R}^{j}:a_{j}^{(1)}>a_{j}^{(2)},\ \ H_{1}^{j}:a_{j}^{(1)}\neq a_{j}^{(2)},\end{split}$
for $j=1,\dots,J$. The $p$-values $\tilde{p}_{j}$ and $1-\tilde{p}_{j}$ correspond to the left and right alternative hypotheses $H_{1,L}^{j}$ and $H_{1,R}^{j}$ respectively. The two-sided $p$-value $p_{j}=2\min(\tilde{p}_{1},1-\tilde{p}_{j})$ corresponds to the two-sided alternative hypothesis $H_{1}^{j}$. The way of computing $\tilde{p}_{j}$ and $p_{j}$ depends on the choice of the univariate test.
One of such methodology to combine individual $p$-values is the well-known Fisher’s method (Fisher, 1932). The Fisher’s combination function is given as:
$P_{L}=-2\sum_{j=1}^{J}\log(\tilde{p}_{j}),\ P_{R}=-2\sum_{j=1}^{J}\log(1-\tilde{p}_{j}),P=-2\sum_{j=1}^{J}\log(p_{j}),$ (4.1)
which correspond to the left, right and two-sided overall alternative hypotheses. $P_{L}$, $P_{R}$ and $P$ all follow $\chi^{2}_{2J}$-distribution. Hence, a final $p$-value is obtained by comparing them with the tail of the $\chi^{2}_{2J}$-distribution.
For the Fisher’s method, $P_{L}$ is used in our illustrative smile examples but not $P$ or $P_{R}$, since we would like to test whether the values of absolute elementary features from Group 1 are smaller than the feature values from Group 2. In other words, we are interested in the left-sided tests. However, the right-sided and two-sided tests could be relevant for some other practical examples.
There is another combination function based on Pearson (1934) which is given as
$Q=\max(P_{L},P_{R}),$ (4.2)
which does not follow a $\chi^{2}$-distribution. Hence, randomization tests and permutation tests are used to estimate the overall $p$-value in two-sample and paired data cases respectively. For the randomization tests, the procedures are given as below:
  1. (a)
    Sample without replacement from the pooled dataset of the two groups for $B$ times.
  2. (b)
    Perform individual two-sample tests on the resampled data and obtain $p$-values $\tilde{p}_{j}^{(b)}$ and $1-\tilde{p}_{j}^{(b)}$, for $j=1,\dots,J$ and $b=1,\dots,B$.
  3. (c)
    Compute the overall $p$-values defined in equation (4.1) and hence $Q$ defined in (4.2).
The procedures for permutation tests are given as the following:
  1. (a)
    Shuffle the subjects randomly, such as change the subjects between experimental group to control group while keeping the paired structure. Repeat this step for $B$ times.
  2. (b)
    Perform individual paired tests to obtain $p$-values $\tilde{p}_{j}^{(b)}$ and $1-\tilde{p}_{j}^{(b)}$, for $j=1,\dots,J$ and $b=1,\dots,B$.
  3. (c)
    Compute the overall $p$-value defined in equation (4.1) and hence $Q$ defined in (4.2).
We use $B=10000$ in both randomization and permutation tests.
The choice between Fisher’s method and Pearson’s method depend on the problem at hand and Owen (2009) provides a review of both methods. In particular, if we suspect there are large deviations from individual null hypotheses, then Pearson’s method has higher power than Fisher’s method; the paper also shows that the Pearson’s method is admissible.
The difference between the feature selection (see Section 3.2) and the Fisher’s and Pearson’s methods is that in the first approach, we record the test statistics resulted in each single test and use an overall test statistics to combine them together. Hence, obtaining a final $p$-value or carrying out feature selection are both possible. In the second approach, we combine the $p$-values obtained from each individual test together via a pre-defined combination function and calculate a final $p$-value. We cannot proceed the feature selection by using this method. The applications of the Fisher’s and Pearson’s methods on the two illustrative smile examples are given in Section 5.2.2, 6.2.2 and 6.3.2 respectively.
5 Analyses of the Cleft Lip Smile Data
5.1 Details